Given: E=I/R and P=IE with the values: P=10 and E=100 What are the values for I and R?

Answer by rothauserc(4718)

Given: E=I/R and P=IE with the values: P=10 and E=100 What are the values for I and R?
P = IE
I = P / E = 10 / 100 = 1 / 10
E = I / R
R = I / E = (1 / 10) / 100 = 1 / 1000
therefore
I = 1/10 and R = 1/1000